Given the following information, determine the cost of the inventory at June 30 using the LIFO perpetual inventory method. June 1 Beginning inventory 15 units at $20 each June 15 Sale of 6 units for $50 each June 29 Purchase 8 units at $25 each The cost of the ending inventory is:

$200

$220

$380

$275

$300
Answer #1

$380

Inventory at the beginning = 15 x $20 = $300

Inventory on 15th June after sale = (15-6) x $20 = $180

Inventory on 30th June = (9 x $20) + (8 x $25) = $380
Answer #2

Given the following information, determine the cost of the inventory at June 30 using the LIFO perpetual inventory method.
 

DateActivitiesUnits Acquired at CostUnits Sold at Retail
Jun. 1Beginning inventory15 units @ $29=$435

Jun. 15Sale




6 units @ $59
Jun. 29Purchase8 units @ $34=$272



The cost of the ending inventory is:

this is a question I don't have an answer to it 

source: Mcgraw Hill connect
answered by: Aaron Dinh